At a University, the faculty members were surveyed and classified according to Designation and number of research publications during the last Academic year. The data are shown in the Table 6. Research Publications Professor Associate Professor Assistant Professor Lecturer 0 – 2 3 6 16 28 3 – 5 2 4 12 8 More than 5 1 1 4 1 Table 6

If a faculty member is selected at random, find the probability that a) The faculty member have 3 publilcations during last year given that he/she is an Assistant Professor b) The faculty member is Lecturer given that he/she do not have any publication
1
Expert's answer
2020-06-14T18:27:47-0400

Given:


Professor Associate Professor Assistant Professor Lecturer TOTAL

0-2 3 6 16 28 53

3-5 2 4 12 8 26

>5 1 1 4 1 7

TOTAL 6 11 32 37 86


Part A)


Formula P(A given B)= P(A and B)/P(B)


P(3 publications given Assistant Professor)=

P(3 publications and Assistant Professor)/P(Assistant Professor)

=12/32=0.375

Part B)

P(Lecturer given 0 publication)=P(Lecturer and 0 publication) /P(0 publication)

=28/53=0.528